OpenStudy (notsmartatall):

determine the constant of variation of each equation (-x^2/y)=3 and y=2x

hero (hero):

To get the constant of variation for the first equation, write it in the form \(\dfrac{y}{x^2} = k\) To get the constant of variation for the second equation, write it in the form \(\dfrac{y}{x} = k\)

OpenStudy (notsmartatall):

@Hero so is the first on y=-3x^2?! lol i dont really get it

hero (hero):

For the first one, divide both sides by -1 to get \(\dfrac{x^2}{y} = -3\) Then multiply both sides by \(y\) Then divide both sides by \(x^2\) Then divide both sides by \(-3\) When you do that, you end up with \(-\dfrac{1}{3} = \dfrac{y}{x^2}\)

hero (hero):

But that is the same as \(\dfrac{y}{x^2} = -\dfrac{1}{3}\) Which means \(k = -\dfrac{1}{3}\)

OpenStudy (notsmartatall):

ohhh ok.. thanks you! what bout the second one?! do i do it exactly the same way too?

hero (hero):

You simply re-write it in the form given above in my first post.

OpenStudy (notsmartatall):

but theres only one number

hero (hero):

For the second one, divide both sides by \(x\) Simple as that